OpenStudy (anonymous):

Find the sum of this geometric series:

OpenStudy (anonymous):

\[\sum_{1}^{\inf} \frac{ (-1)^(n-1) }{ 5^n }\]

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

Sorry this is (-1) ^ (n-1)

OpenStudy (anonymous):

lol, this look like Maclaurin Series, you need to force it into a common equation

OpenStudy (anonymous):

Its geometric, the question says it.

OpenStudy (anonymous):

I lied, 1/6 should be the answer

OpenStudy (anonymous):

the common ratio is \(-\frac{1}{5}\) and the first term is \(\frac{1}{5}\) so it is \[\frac{\frac{1}{5}}{1+\frac{1}{5}}=\frac{1}{6}\]

OpenStudy (anonymous):

Oh okay thanks a lot guys! I thought the common ratio was 1/5 not -1/5